OpenStudy (daisyduck04):

will fan and medal plz help me Stephanie's rectangular garden was 12 feet long and had a perimeter of 40 feet. She made her garden 6 feet longer. How could she calculate the area of her new garden?

OpenStudy (daisyduck04):

A. She could calculate that the length of her new garden is 18 feet. The width of this garden would be 2 feet if the perimeter is 40 feet. She would multiply 18 × 2 to calculate the area. B. She could calculate that the dimensions of her original garden were 12 feet by 8 feet. The length of the new garden would be 18 feet. She would multiply 18 × 8 to calculate the area. C. She could calculate that the width of her original garden is 28 feet. The length of the new garden would be 18 feet. She would multiply 28 × 18 to calculate the area.

OpenStudy (daisyduck04):

@help?!?!?! @MysticalFlames

OpenStudy (anonymous):

is this algebra you use a2+b2=c2

OpenStudy (anonymous):

so then do want the long answer or short

OpenStudy (daisyduck04):

no not algebra

OpenStudy (daisyduck04):

i think its perimeter

OpenStudy (anonymous):

ok then it would be a becuse remeber perimeter is when you add all sides together so if you add 12+6=18 right

OpenStudy (anonymous):

do you get it

OpenStudy (anonymous):

or want me to explain

OpenStudy (daisyduck04):

sorry my laptop died so i went on my computer

OpenStudy (daisyduck04):

gave you a medal and fanned you

OpenStudy (daisyduck04):

@Cruznatalie150

OpenStudy (anonymous):

it is 40-12-12 = 2b 16=2b 8=b = width new garden (12+6)*8= 144 ft square

OpenStudy (daisyduck04):

um thanks

OpenStudy (anonymous):

welcome.